cashews sell for $2.25/kg and peanuts for $1.75/kg. How much of each type must be mixed to make 20kg of a mixture that will sell for $41. Thank you!
Answer by ewatrrr(24785)

2.25x + 1.75(20kg - x) = 2.05(20kg) Note: $41/20 = $2.05
x = .30(20kg)/.5, amount at $2.25